Factor the polynomials using whatever strategy seem appropriate. State which factoring methods to use and then demonstrate them methods on your problems, explaining the process while you go. Discuss any challenges those particular polynomials posed for the factoring.
1st One:
t²+4z²
2nd One:
y³+y²-4y-4

Answer by stanbon(75887)

Factor the polynomials using whatever strategy seem appropriate. State which factoring methods to use and then demonstrate the methods on your problems, explaining the process while you go. Discuss any challenges those particular polynomials posed for the factoring.
1st One:
t²+4z² = (t-2zi)(t+2zi)
-----------------------------
2nd One:
y³+y²-4y-4
Rearrange:
y^2(y+1)-4(y+1)
--------
(y+1)(y^2-4)
--------
= (y+1)(y-2)(y+2)
=======================
Cheers,
Stan H.
============